Find work done in stretching a light elastic string to double its length???

Grade:12th pass

1 Answers

Dibyakanti Kumar
6 years ago
Integrate F*dx =k*x *dx where k is the elasticity constant . Integrating from l to 2l
(where l is initial length) we get 0.5*k*3*l^2 =1.5*k*l^2 which is the work done.